OpenStudy (joshoyen):

Solve each equation by factoring. 1. m2-36=0 2. v2+7v-8=0 3. v2-9v=-8

OpenStudy (dayakar):

1) \[a^{2}-b^{2} =(a+b)(a-b)\] use this factorise

OpenStudy (dayakar):

can u do it

OpenStudy (joshoyen):

Hmm, im still confused

OpenStudy (dayakar):

\[m^{2}-6^{2}=0\]

OpenStudy (joshoyen):

ohh

OpenStudy (dayakar):

(m+6)(m-6)=0

OpenStudy (joshoyen):

alright thanks!

OpenStudy (dayakar):

what are the m values?

OpenStudy (joshoyen):

(-6,6) right?

OpenStudy (dayakar):

very good

OpenStudy (joshoyen):

Ok thank you very much

OpenStudy (dayakar):

2) split the middle term

OpenStudy (joshoyen):

v2-1v+8v-8=0 ?

OpenStudy (dayakar):

excellent

OpenStudy (joshoyen):

alright im just a bit confused on the last steps

OpenStudy (dayakar):

v(v-1)+8(v-1)=0

OpenStudy (dayakar):

r u clear

OpenStudy (joshoyen):

yep , thanks!

OpenStudy (joshoyen):

it would be (1, -8) ?

OpenStudy (joshoyen):

wait how would i do x2+x-56=0

OpenStudy (dayakar):

can u do remaining part

OpenStudy (joshoyen):

yep, i'm all good now, thanks alot!

OpenStudy (dayakar):

ok

OpenStudy (joshoyen):

when wait you split the equation how do you know which numbers go where, so like for p2+9p+18, it could be p2+6p+3p+18 or p2+3p+6p+18

OpenStudy (dayakar):

u can write either way u will get same answer for both

OpenStudy (dayakar):

\[3n^{2}-6n+n-2=0\]

OpenStudy (dayakar):

is it clear

OpenStudy (joshoyen):

really? but say for instance the answer -6,- 3 on my answer sheet in my book it says -3, -6

OpenStudy (dayakar):

it makes no difference p= -3 or p= -6 p= -6 or p= -3 is there any difference

OpenStudy (joshoyen):

ohhh, okay wait nvm I see, i kept thinking it was (x, y) format